Determine the GMR the unconventional stranded conductors shown below 20 radius of each conductor is r 21 21
Answer #1

se 2r Radius of stranded conductor is r. 19GMR (a) (23) (2(SE) Here d = 8+27+8 = 40. GMR = 16(elu (ur) (88) (or) (8522 = 160.7788 2896-31744 = 19 (2255.65%. 86 = (2255.